the age of the father is twice the sum of the ages of his two children,after 20yearshis age wil be equal to the sum of the ages of his children,find the age of the father
Answer by jorel1380(3719)

F=2K
F+20=K+40
F=K+20
2K=K+20
K=20
F=40
The father is 40. His two kids together are 20..